# Customize-It-3D: High-Quality 3D Creation from A Single Image Using Subject-Specific Knowledge Prior

### v1

@@ 3.3 Refine Stage: Neural Texture Enhancement
Figure 7: Qualitative comparison on image-to-3D generation. We compare Customize-It-3D to RealFusion [38], Make-it-3D [59] and Magic123 [46] for creating 3D objects from a single unposed image (the leftmost column).
Deferred point cloud rendering and training. After initializing the geometry and texture of the point cloud $P$, a focus is placed on enhancing the texture of the point cloud in novel views with subject-specific prior. Specifically, we assign a 19-dimensional descriptor $F$ to each point in the novel view $\mathbf{v}$, where the first three dimensions are colors initialized from the above process. During the rendering of the point cloud, we employ a deferred rendering scheme [59], rendering the point cloud K times at different resolutions and concatenating the resulting K feature maps into the final image using a jointly optimized U-Net network denoted as $\mathcal{R}_{\theta}$. The refine stage’s loss is similar to that in Sec. 3.2, with the addition of a regularization term penalizing excessive differences in texture between the refined and initialized states.
4 Experiments
4.1 Implementation Details
@@
[…]
Table 1: Quantitative comparison on RealFusion15. We compute LPIPS and PSNR under the reference view, and CLIP under novel views.
[…]
NeRF and point cloud rendering. We employ multi-scale hash encoding from Instant-NGP [41] and maintain an occupancy grid to skip empty space for improving rendering efficiency. Additionally, we incorporate shading augmentations, such as Lambertian shading similar to [45]. However, during the first 1000 iterations of the second phase of progressive training, we utilize normal shading to emphasize the learning of better geometry. For the deferred rendering strategy of point clouds, we utilize a 2D U-Net architecture with gated convolutions [71].
4.2 Comparisons with the State of the Arts
@@
Baselines. We compare our approach against recent state-of-the-art methods: RealFusion [38], Make-It-3D [59], and Magic123  [46] (using Zero123 XL checkpoint  [34] as ours for fair comparison). We compare these methods using the Realfusion dataset [38] and a customized dataset we created. The Realfusion dataset includes many natural images, while our customized dataset comprises real images and images generated by Stable Diffusion XL [44]. We evaluate all the baseline methods using their official code.
Figure 8: The effect of Multi-modal DreamBooth, showing a consistent improvement in terms of both geometry and texture.
Figure 9: Single RGB versus Multi-modal DreamBooth, showing a clear improvement particularly in geometry.
@@
Qualitative comparison. We present a comprehensive collection of qualitative results in Figure 7. RealFusion often generates flat 3D results with colors and shapes that diverge significantly from the input image. Make-it-3D exhibits competitive texture quality but suffers from an issue known as long geometry in side views, particularly noticeable in the reconstruction of objects such as chairs. Magic123 produces visually plausible structures but grapples with a notable issue of multi-face, as it tends to replicate the reference texture in the back view.. In contrast, our approach can reasonably hallucinate the texture details and geometry for novel views that even deviate significantly from the reference image, which greatly improves the fidelity and consistency in creating 3D models.
@@
Quantitative comparison. We quantitatively compare with the baselines in Table 1. We use metrics following [59, 46] with PSNR, LPIPS [74], and CLIP-similarity [47]. As shown in the table, Customize-It-3D achieves Top-1 performance across all the metrics, showing that our model is able to generate 3D objects with better 3D consistency.
5 Ablations and Analysis
With or without multi-modal DreamBooth. We first ablate the effect of using the proposed multi-modal Dreambooth in Figure 8. We observe a consistent improvement in terms of both texture and geometry. Employing the suggested multi-modal subject-specific diffusion model prior results in higher-quality 3D content, yielding more compelling visuals with enhanced 3D consistency.
Single RGB versus multi-modal. We experiment a baseline that employs conventional DreamBooth finetuned from a single RGB reference to further investigate the impact of our shading mode-aware multi-modal guidance. The difference shown in Figure 9 demonstrate a clear improvement, whereas the baseline suffers from multi-face problem.
Figure 10: Illustrating high-fidelity text-to-3D generation.
Figure 11: Illustrating high-fidelity 3D creation for real scenes.
6 Applications
High-quality text-to-3D generation. To achieve high-quality text-to-3D creation, we initially utilize a T2I diffusion model to convert the text prompt into a reference image and then pass it into our image-based 3D creation method. We show the results in Figure 10. Customize-It-3D demonstrates remarkable quality in text-to-3D generation.
Real scene modeling. We extend our model evaluation to real-world scenarios by inputting arbitrary general images capturing complex, authentic scenes into our framework. Figure 11 presents several visual examples, verifying the efficacy of our method in modeling real scenes with a high degree of capability and fidelity.
7 Conclusion
We have presented Customize-It-3D, a novel two-stage approach for image-to-3D generation. Leveraging the established customized knowledge prior, we enhances NeRF optimization for improved geometry and refines texture, resulting in superior alignment of 3D content with the subject. Customize-It-3D demonstrates versatility in handling general objects and empowers captivating applications.
Supplementary Material
Figure 12: Additional comparison with baseline: RealFusion [38], Make-It-3D [59], and Magic123  [46]. The first column is the reference image.
A Additional Implementation Details
A.1 Coarse stage
Multi-modal Dreambooth finetuning and usage. We utilized the stable diffusion V2.0 model [51] in our experiments. We employed multi-modal images related to the reference image and conducted a four-step fine-tuning process on the diffusion model. The sequence and number of iterations for each training step are as follows: 25 iterations for mask image training, 50 iterations for depth image training, 50 iterations for normal image training, and 75 iterations for RGB image training, totaling 200 iterations. During fine-tuning, prompts were specified as follows: “a foreground mask / depth map / normal map / RGB photo of sks [class_name].” When using the finetuned model for training in the coarse stage, we modified the prompts based on shading mode. For normal shading, we append the modifier “normal map” to the text prompt, while for other shading strategies, we add the modifier “RGB photo”. We observed that this approach facilitated the better utilization of subject-specific knowledge, leading to improved results.
Optimizing the pipeline. We employ the Adam optimizer [67] with a learning rate of 0.005 and no weight decay for 7000 iterations. Simultaneously, during the coarse stage, we implement progressive training, where the first 2000 iterations exclusively trained within the range of ±$\pm 45^{\circ}$ from the reference view, and the subsequent 5000 iterations covered the entire $360^{\circ}$ around the object. In the refine stage, we conducted 3000 iterations to optimize the texture.
Camera setting. For the camera sampling method, we randomly sample novel views with a probability of 75%, while with a 25% probability, we sample a reference view following the approach in [45]. The reference view is assumed to be a frontal view, i.e., a polar angle of $90^{\circ}$ and an azimuth angle of $0^{\circ}$. It is further assumed that the camera is positioned at a distance of 1 meter from the origin of the coordinate system, and the field of view (FOV) is set to $60^{\circ}$.
A.2 Refine stage
Post-mask processing. We also perform post-mask-processing, wherein we apply certain morphological operations to both SAM-generated and NeRF-generated masks, i.e., taking the union of the two masks to complement fine structures lacking in the initial mask. Finally, we have $M_{pseudo}=\hat{M}\cup\hat{M}_{SAM}$, where $\hat{M}$ and $\hat{M}_{sam}$ are the rendering and SAM-generated mask after post-processing under novel view.
Point cloud building. The goal of our point cloud building strategy is to avoid introducing points with color conflicts that overlap with existing point cloud $p^{m}_{ref}$, which is build from reference view. We use a reference mask $M^{ref}_{pseudo}$ to perform color mapping on $p^{m}_{ref}$, obtaining the front-facing perspective of the new point cloud $p_{ref}$. For all remaining regions under novel view, we project the just obtained point cloud $p_{ref}$ onto each novel view $v_{i}$ to create a new mask $m^{re}_{i}$. We then take the set difference between the novel view mask $m^{i}_{pseudo}$ and the re-projected mask $m^{re}_{i}$ as the mask for the remaining points $P_{left}=P^{m}-p^{m}_{ref}$. Finally we will have a clean point cloud $P=\{p_{ref},p_{1},...,p_{n}\}$ which ensures color mapping is done without introducing conflicts.
Point cloud rendering. In order to perform better point cloud rendering, each point’s has a 19-dimensional descriptor, with the initial three dimensions initialized with RGB colors and the rest initialized randomly [59]. Deferred rendering is conducted for a total of $K=3$ times, with resolutions denoted as $[H/2^{i},W/2^{i}]$, where $i\in[0,K)$ and ultimately, the learned U-Net network concatenates these three feature maps. And the U-Net architecture contains 3 down-sampling and up-sampling layers with gated convolutions[71]. The rendering resolution is set to $128\times 128$ in the coarse stage for NeRF and $800\times 800$ in the refine stage for the point cloud, respectively.
Figure 13: Additional results of an ablation study on whether to use the multi-modal dreambooth. The first column is the reference image.
Figure 14: Additional ablation results. The first column is the reference image. The second column displays the results obtained using our method. The third column showcases the outcomes when substituting the multi-modal dreambooth with the traditional dreambooth. The fourth column exhibits the results when shading mode guidance is omitted during the NeRF training process.
B Additional Ablation Study Results
As mentioned in the paper, current methods for text/image-to-3D rely on a generic diffusion model with strong generalization capabilities, leveraging implicit 2D knowledge to guide the generation of novel views. However, due to its overly rich imagination, this model may generate 3D objects that do not align with the subject identity. To address this limitation, we fine-tune the diffusion model to learn subject-oriented knowledge. We conduct a series of experiments to demonstrate the effectiveness of our subject-specific and multi-modal diffusion model. We first ablate the effect of using the proposed multi-modal Dreambooth, as illustrated in Figure 13. The results indicate that without the guidance of our multi-modal Dreambooth, the coarse stage produces textures and geometry that deviate from the subject identity, thereby impacting the initialization of point cloud geometry and textures in the refine stage, as exemplified by the patterns on the fox’s neck and the rabbit’s ears.
Furthermore, we compare the results obtained using the traditional dreambooth (trained with a single RGB image) with our multi-modal dreambooth (trained with multiple multi-modal images). We also compare the results of the NeRF training process with and without shading mode guidance. Both sets of experiments are depicted in Figure 14. The results show that a simple attempt using only dreambooth leads to overfitting and mode collapse. Additionally, we observed that using shading mode guidance better leverages our multi-modal dreambooth, resulting in more convergent and 3D-consistent outcomes.
[…]
Table 2: Quantitative comparison on our dataset. We compute LPIPS, PSNR and CLIP-similarity.
C Additional Comparisons with Baseline
In this section, we show more results comparing our method with the baseline: RealFusion [38], Make-It-3D [59], and Magic123  [46] (using Zero123 XL checkpoint [34]), as shown in the Figure 12. Also we do quantitative comparison on our own dataset as shown in the Table 2, and our method has the best results. Results show that our method exhibits increased fidelity to the input reference image in both geometry and texture, thanks to the utilization of subject-specific knowledge priors. This incorporation makes the training process more aligned with the subject’s identity. Particularly in terms of texture, as exemplified in the second row with the silver cat in the Figure 12, our method generates back textures with a more pronounced metallic sheen, consistent with the characteristics of the front texture. Similarly, as seen in the fifth row with the doll in the Figure 12, our method generates doll textures on the back and sides, including patterns on the skirt and hair color, that better align with the subject’s identity. Therefore, our method is capable of generating textures that are more consistent with the reference view.
D Additional Results
In this section, we show more results using our method. Figure 15 and Figure 16 demonstrates the ability of our method to generate high-quality 3D objects from a single image.
E Limitation
Our method has several limitations. Firstly, because our approach relies on using pretrained models as depth estimator [50] and normal estimator [12] [23] to estimate the depth map and normal map of the reference image. So any errors in these modules will affect our method and impact the overall generation quality. Also, our method, relying on generative prior to imagine the 3D content from the single reference viewpoint, inevitably introduces geometry ambiguity [45], leading to issues such as the Janus problem or over-flat geometry.
Figure 15: Additional results by Customize-It-3D. The first column is the reference image.
Figure 16: Additional results by Customize-It-3D. The first column is the reference image.

### v3

@@ III-B Coarse Stage: Image to 3D Reconstruction
Figure 3: Qualitative comparison on image-to-3D generation. We compare Customize-It-3D to RealFusion [40], Make-it-3D [64] , Magic123 [49] and DreamGaussian [63] for creating 3D objects from a single unposed image (the leftmost column).
Figure 4: Comparison with Magic123 in coarse stage.
III-C Refine Stage: Neural Texture Enhancement
In the coarse stage, due to the computationally intensive nature of NeRF optimization, we only perform optimization at low resolutions. Moreover, due to the inherent limitations of NeRF, such as the tendency to produce high-frequency artifacts [48], we can only obtain a low-resolution, low-texture-quality 3D model. As a result, we are motivated to explore alternative representations for NeRF, particularly focusing on transforming it into an explicit point cloud. Point clouds have the advantage of allowing direct projection of images, preserving high-quality frontal texture details, and offering opportunities for personalization and enhancement of the projected images, fulfilling the needs of high-quality customized 3D models. In summary, our goal in the refine stage is to retain the geometric structure of the coarse stage, generate a dense point cloud, and optimize the textures not visible in the reference view $\mathbf{v}_{ref}$, as we directly project the frontal reference image $\mathbf{x}$ onto the point cloud. We elaborate several key designs to facilitate this transformation.
Point cloud building. Past work has focused on constructing a point cloud from multi-view RGBD images under NeRF rendering [64]. However, due to noise in RGBD images and inaccurate depth estimation, the generated point cloud is very noisy , significantly impairing the 3D geometry in refine stage. To address this, we export the NeRF in coarse stage as a mesh, utilizing mesh regularizations and perform Poisson sampling on the mesh to obtain a dense point cloud $P^{m}$. Figure 2 shows this difference. We opt to a point cloud because it allows personalized diffusion model-based texture enhancement for each viewpoint, resulting in a higher-quality point cloud that better aligns with our customized 3D object generation needs.
Texture projection. However, since NeRF-rendered images from different viewpoints may have overlapping regions, a 3D point may be assigned different colors under different viewpoints [71], leading to texture conflicts. Therefore, we propose an iterative strategy for constructing texture point cloud $P$ from mesh sampled point cloud $P^{m}$. Initially, we trust all points $P^{m}_{ref}$ under reference view $\mathbf{v}_{ref}$ and perform color mapping on $P^{m}_{ref}$, obtaining the front-facing perspective of the new point cloud $P_{ref}$. For novel view projection, we aim to avoid introducing points with color conflicts that overlap with $P_{ref}$. Hence, we proceed to reproject the recently obtained point cloud $P_{ref}$ onto each novel view $\mathbf{v}_{i}$ to generate a new mask denoted as $M^{ref\_proj}_{\mathbf{v}_{i}}$. We then project textures to the points under this novel view only when these points belong to the set difference between the novel view mask $M_{\mathbf{v}_{i}}$ and the reprojected mask $M^{ref\_proj}_{\mathbf{v}_{i}}$. Finally we have a clean point cloud $P=\{p_{ref},p_{1},...,p_{n}\}$ which ensures color mapping is done without introducing conflicts.
Texture and mask enhancement. The texture quality obtained in the coarse stage is suboptimal. Therefore, prior to projecting the reference image $\mathbf{x}$ from the reference view and rendered RGB images $\mathcal{X}=\{\mathbf{x}_{\mathbf{v}_{1}},\mathbf{x}_{\mathbf{v}_{2}},...,\mathbf{x}_{\mathbf{v}_{n}}\}$ under novel views onto the point cloud geometry, we perform a texture enhancement process on all rendered images within $\mathcal{X}$. Specifically, we add noise to the rendered images and leverage the previously fine-tuned multi-modal personalized diffusion model to denoise into a set of pseudo images $\mathcal{X}^{pseudo}$ that possess higher texture quality as well as multi-view consistency. Simultaneously, due to the issues in the coarse stage of NeRF, the generated masks =$\mathcal{M}=\{M_{\mathbf{v}_{1}},M_{\mathbf{v}_{2}},...,M_{\mathbf{v}_{n}}\}$ by NeRF are not satisfactory. Fine details (for example long and thin parts) in the object structure may be compromised, leading to defects in point cloud texture and geometry initialization due to mask inadequacies. Therefore, we adopt Segment Anything Model (SAM) [28] to generate more refined and accurate masks $\mathcal{M}^{pseudo}$. However, due to the insufficient resolution of RGB images and the noise introduced by SAM, we assess the quality of the generated mask by relying on SAM’s score. Only when the quality surpasses a certain threshold do we consider using this mask; otherwise, we adhere to utilizing masks generated by NeRF.
Figure 5: The effect of multi-modal DreamBooth.
Figure 6: Our proposed method multi-modal DreamBooth with shading mode-aware guidance shows a clear improvement. All results are obtained under novel view in the coarse stage.
(a) Mask details are better preserved.
(b) Texture details are enhanced by Multi-modal DreamBooth.
Figure 7: Results for mask processing and texture enhancement.
Deferred point cloud rendering and training. Here we adopt a process akin to that described in Make-it-3D [64] for point cloud rendering and training, but we incorporate our subject-specific prior. It is worth to mention that the point cloud building and texture projection is different, which we add an illustration figure in the appendix for better clarity.
Dataset Metrics RealFusion Make-it-3D Magic123 DreamGS Ours
RealFusion CLIP↑ 0.73 0.83 0.82 0.77 0.90
PSNR↑ 16.87 20.01 19.50 18.94 20.50
LPIPS↓ 0.193 0.119 0.100 0.111 0.094
Test-data CLIP↑ 0.70 0.79 0.86 0.79 0.88
PSNR↑ 16.41 19.31 20.10 19.80 20.30
LPIPS↓ 0.197 0.121 0.098 0.103 0.096
TABLE I: Comparison of different methods on RealFusion15 and Test-data. We compute LPIPS and PSNR under the reference view, and CLIP- Score under novel views.
@@ III-C Refine Stage: Neural Texture Enhancement
TABLE II: The quantitative ablation study result.
IV Experiments
IV-A Comparisons with the State of the Arts
@@
Baselines. We compare our approach against recent state-of-the-art methods: RealFusion [40], Make-It-3D [64], Magic123  [49] and DreamGaussian [63]. We compare these methods using the Realfusion dataset [40] and a customized dataset we created. The Realfusion dataset includes many natural images, while our customized dataset comprises real images and images generated by Stable Diffusion XL [47].
@@
Qualitative comparison. We present a comprehensive collection of qualitative results in Figure 3. RealFusion often generates flat 3D results with colors and shapes that diverge significantly from the input image. Make-it-3D exhibits competitive texture quality but suffers from an issue known as long geometry in side views, particularly noticeable in the reconstruction of objects such as chairs. Magic123 produces visually plausible structures but grapples with a notable issue of multi-face, as it tends to replicate the reference texture in the back view. The geometry generated by DreamGaussian seems to be highly disordered. Under novel views, the newly generated textures are also of poor quality, lacking in detail.
To further demonstrate the effectiveness of the proposed multi-modal DreamBooth, we present results from the coarse stage without any components in the refine stage in Figure 4 compared to Magic123[49]. It can be seen that our results achieve superior results in both geometry and texture, e.g., ours correctly generate the back view of anime girl while Magic123 suffers from multi-face issue.
@@
Quantitative comparison. We quantitatively compare with the baselines in Table I. We use metrics following [64, 49] with PSNR, LPIPS [78], and CLIP-similarity [50]. As shown in the table, we achieves best performance.
V Conclusions and Discussions
Subject-specific prior: multi-modal DreamBooth. We first ablate the effect of using the proposed multi-modal DreamBooth in Figure 5. We observe a consistent improvement in terms of both texture and geometry. Take banana for instance, without multi-modal DreamBooth, the method fails to reconstruct slender structures already visible in the input banana image, which also appears to have unsuccessful reconstruction in Make-it-3D such as the armchair leg shown in Fig 3. The reason is inconsistency stemmed from SDS loss (novel view) and reconstruction loss (reference view). To achieve a balance between these , Magic123 addresses this with additional 3D prior[34], and we alleviates it through multi-modal DreamBooth. This example indeed demonstrate the efficacy of our proposed multi-modal DreamBooth. Also, we conduct a quantitative study on our test benchmark, validating the efficacy of our approach in Table II.
Shading-mode-aware guidance. We ablate shading-mode guidance and found that with it, our training in the coarse stage converged more easily, showing better performance in both geometry and texture, as illustrated in second column and fourth column of Figure 6. This fully demonstrates that shading-mode guidance effectively utilizes the subject’s multi-modal information to provide more robust supervision.
Texture and mask enhancement. Here we ablate the proposed texture and mask enhancement procedures. Figure 7(b) shows that the texture details can be significantly improved by utilizing our multi-modal DreamBooth model. As shown in Figure 7(a), our method can better retain the geometry.
VI Acknowledgement
This work was supported by the National Natural Science Foundation of China (62476011).